$SPAD/src/input d001amf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
e:EXPR FLOAT:=1/((X+1)*sqrt(X))
f:ASP1(F):=retract e
a:SF:=0.0
epsabs:SF:=0.0
epsrel:SF:=1.0e-4
liw:=200
lw:=4*liw
result:=d01amf(a,1,epsabs,epsrel,lw,liw,-1,f)

———-

References^{[1] nothing}

4